OpenStudy (anonymous):

What is the probability that two people selected at random have the same birthday? Ignore leap years. 1/365 2/365 1/133,225

OpenStudy (anonymous):

is it three?

OpenStudy (ash2326):

@sdk First person could have his birthday on any of the 365 days \[\frac{365}{365}\] The probability of other having his/her birthday on the same day \[\frac{365}{365} \times \frac{1}{365}\] Can you simplify this?